OpenStudy (anonymous):

PLEEEEEASE HELP ME FINISH THIS!! I finished task 1 already, here is what I did: Create expressions to represent the dimensions of the pool. a.Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable. b.Determine the polynomial expression to represent the area of land space that the pool will cover. c.Using the area and the depth expressions, determine the polynomial expression that is used to represent the volume. d.Classify the volume polynomial by degree and number of terms.

OpenStudy (anonymous):

a. Width= x Length= (3*x)+3 Depth= (6*x) -1 b.Area of Land Space Area= L*W (3x+3)*x 3x^2+3x c. Volume= A*D (3x+3)*x*(6x-1) 18x^3+15x^2-3x d.The number of terms are three and the highest degree is three

OpenStudy (anonymous):

THIS IS WHAT I NEED HELP WITH: Task 2 Bernard is realizing that his dimensions are not fitting the way that he wants it to because he forgot about the deck. The deck is going to be a rectangular space that borders the pool. Also, for the pool to fit just right, there must be a special product represented when either the area or volume expressions are determined. He wants you to design your own rectangular pool using polynomial expressions, including the deck that borders the pool. a.Include a diagram or sketch of your pool and deck. Your diagram must have all dimensions represented with polynomials and ensure that a special product will be used. b.Determine the polynomial expression to represent the area of land space that the pool will cover. c.Determine the area of the deck that borders the pool based on your dimensions. Be sure to include all of your work and calculations to receive full credit

OpenStudy (phi):

this sounds like some kind of torture.

OpenStudy (phi):

did you make a sketch?

OpenStudy (phi):

also, what do they mean by "special product" in there must be a special product represented when either the area or volume expressions are determined. ?

OpenStudy (anonymous):

@iambatman @ninjasandtigers @uri @skullpatrol @SolomonZelman @camerondoherty SOMEONE HELP! it doesn't have to be 100% right. i just have to show effort

OpenStudy (solomonzelman):

Why are you tagging me, if PHI is replying to your question ?

OpenStudy (camerondoherty):

^

OpenStudy (anonymous):

he isn't, he's as lost as i am XD

OpenStudy (anonymous):

:(

OpenStudy (solomonzelman):

I read very long problems during the winter season, but I don't like reading them during summer. I am a lazy A__─Hole -:( Phi isn't lost ... I've seen do harder roller coaster, BUT the instructions must be clear before you do anything, and that's what he is saying.

OpenStudy (anonymous):

yeah, i got it. no one is going to help me with this, its too long. i wouldn't either. I'll just take whatever the teacher can give me for what I've done @SolomonZelman

OpenStudy (solomonzelman):

Phi is willing to do it.. but he wants clear instructions.

OpenStudy (camerondoherty):

<aybe to make it less confusing... ask a new question without task 1 and only task 2...

OpenStudy (camerondoherty):

*maybe

OpenStudy (anonymous):

@camerondoherty but don't you need the info from task 1 to do task 2? or no?

OpenStudy (camerondoherty):

Oh... I didnt notice... But you dont need to put the whole question and such... Just put the old dimensions of the pool and all of the nessesary information that is needed for the next task

OpenStudy (anonymous):

okay, thanks!

OpenStudy (camerondoherty):

Making it look nicer and less complicated helps. The more organized you are the easier it s to solve the problem c;